OpenStudy (anonymous):

quadratic equation by completing the square. check both roots . equation below

OpenStudy (anonymous):

\[x^2-6x=7\]

OpenStudy (anonymous):

please explain each step

OpenStudy (nottim):

well, first subtract 7 from both sides.

OpenStudy (anonymous):

take half of 6 get 3 write \[(x-3)^2=7+...\] then think "3 squared is 9" write \[(x-3)^2=7+9=16\] take the root, don't forget the \(\pm\) get \[x-3=\pm\sqrt{16}\] \[x-3=\pm4\]

OpenStudy (anonymous):

then solve \[x-3=4\iff x=7\] or \[x-3=-4\iff x=-1\]

OpenStudy (anonymous):

all the steps are there

OpenStudy (anonymous):

wow i think i get it let me try it on paper

OpenStudy (anonymous):

hmm thats not right

OpenStudy (anonymous):

wait my mistake you got half right but there is supposed to be one more answer

OpenStudy (nottim):

no...satellite gave enough I think. There's 2 values for x in his answer. I dont think there's a third one is a trinomial like this.

OpenStudy (anonymous):

you are right my bad again i didnt see or